# Displacement memory for flyby

### v4

@@ V Higher-derivative profiles
Similar results can be obtained for the higher-derivative profiles proposed in GibbHaw71 ; LongMemory ; Benin .
@@
This curious “half-sidedness” is explained as follows. Assuming that $k>0$, DM is obtained for $X^{2}$ by fine-tuning the amplitude, $k=k_{crit}^{(2)}$. This does not work for $X^{1}$, though, unless $X^{1}\equiv 0$, as confirmed by FIG.13.
Figure 13: Fine-tuning one of the sectors for DM yields VM for the other component, unless the latter is identically zero.
We could have proceed also with the opposite casting and fine-tune the amplitude for $X^{1}$ to get $k=k_{crit}^{(1)}$. But the profiles have opposite signs, therefore
@@
Models with higher-order derivatives of the Gaussian follow a similar pattern. For the 4th order derivative profile, for example, we would get again half-DM effect, etc.
We now explain the reason by extending our observation related to the behaviour under reversing the sign of $U$, $U\to-U\,$ in (IV.21), We start with an even seed profile,
${{\mathcal{A}}}^{0}(-U)={{\mathcal{A}}}^{0}(U)\,,$ (V.5)
@@
The sign-reversal (IV.21) leaves invariant the 2nd-order $U$-derivative terms in (I.3) but has an $\ell$-dependent effect on the equations of motion :
@@
implying that the two transverse components in (I.3a)-(I.3b) are merely interchanged under (IV.21) (as we noticed for flyby in sect.III). Thus they satisfy the same pair of equations up to permutation. Therefore the DM condition holds either for both or for none of them. When it does, then “full DM” is obtained for both coordinates as illustrated in FIGs. 8-9 for $\ell=1$, and in FIG.15 for $\ell=3$. $U$-reversal symmetry (IV.21) allows us to generalize our observations in sect.IV.
@@
Each of the transverse equations (I.3a) and (I.3b) is left invariant under (IV.21)
@@ VI Longitudinal motion
$\mathfrak{m}=M\,.$ (VI.4)
Then setting $Z=V+\frac{1}{2}U$ we get,
$Z(U)=V_{0}=\mathop{\rm const.}\nolimits$ (VI.5)
i.e.,  we have DM also in the longitudinal $Z$-direction, see FIG.#(14) of DM-1 .
@@ VII Conclusion
Test particles exhibit, generically, the velocity effect (VM) Ehlers ; Sou73 ; AiBalasin  : when hit by a sandwich gravitational wave , the particles fly apart with diverging constant, non-zero velocity ShortMemory ; LongMemory .
Zel’dovich and Polnarev suggested instead that flyby would approximately generate pure displacement (DM) with vanishing relative velocity ZelPol . Our main result is to confirm and indeed refine the statement of Zel’dovich and Polnarev. Both numerical and analytical evidence show that a judicious choice of the wave parameters yields pure displacement DM-1 .
We emphasise also that the very notion of “wave zone” is problematic. It is, as indicated already by Zeldovich and Polnarev ZelPol , only approximate. How could one define where a Gaussian “ends" ? The answer would depend on the precision. It means, intuitively, “very small”. But if we look at it with a magnifying glass of (say) 100, a trajectory which looked to be constant will become …oscillating. So the only phrase that we can say rigorously: “the amplitude is smaller than epsilon - and the region where this is so depends on epsilon. We have however a semi-analytical argument: the critical amplitudes listed (IV.16) are the roots of (IV.11) (shown in FIG.11) that we could determine numerically to some accuracy.
Another important clarification about “quantization” : our theory is purely classical, we do not consider a quantum theory : we use the word in the sense of Old Quantum Mechanics highlighted by the names of Planck, Bohr, Sommerfeld, etc, who found that certain physical quantities as energy, angular momentum, etc should be “quantized” meaning that they should take discrete values. Remember that this happened before present-day Quantum Mechanics was born.
What is astonishing, though, is that the analogy between classical mechanics/relativity theory and QM goes beyond just “taking discrete values”. As noticed in DM-1 , our DM trajectories correspond in fact to solutions of the time-indendent Schrödinger equation with growth conditions at infinity. However, no normalizability of the type $\int|\Psi|^{2}<\infty$ is required; on the contrary, we are looking for zero “energy", non-normalizable solutions, which are not considered in quantum mechanics.
In this paper we study gravitational waves with linearly polarized “plus” profile. Those with “cross” profile are obtained by a rigid rotation and lead to analogous results. Combining the two polarizations is under current investigation.
More generally, we consider gravitational waves whose profile is a derivative of order =0, $\ell=0,\,1,\ \dots$ of a seed profile ${\mathcal{A}}^{0}$, (V.6). The latter can be, for example, a Gaussian GibbHaw71 or its analytical Doppelgänger, the Pöschl-Teller potential PTeller ; Chakra ; DM-1 . Other promising candidates (not studied in the paper) could be Rosen-Morse type Scarf or inverse-power AndrPrenc ; Ilderton ; Benin profiles. Analytically solvable models can be built also from square-profile Chakraborty:2022qvv ; Benin which have a well-defined wave zone.
@@
In this paper we limit our attention at particles at rest before the wave arrives. It is legitimate to wonder, though, what would happen for nonzero initial velocity. In the impulsive case AiBalasin ; impulsive ; Steinbauer we found merely diverging straight trajectories and no DM, see FIG.#7 of impulsive .
Preliminary versions of this research were presented by PAH at the Wigner Institute in Budapest (April 5 2024), at the Workshop Carrollian Physics and Holography-CDFG-2024 organised at the Schrödinger Institute in Vienna (April 17 2024) and at the conference “Conformal anomalies: theory and applications 2024, https://indico.math.cnrs.fr/event/10718/ Tours (May 7 2024). P-M Zhang presented related results in Benin . Similar ideas were put forward also in Bondi57 ; Sou73 ; AiBalasin ; GriPol . Recent results on the Memory Effect and the DM or VM puzzle include refs. Chakraborty:2022qvv ; Mitman ; HarteOancea ; Arpan ; DeyKar .
@@ Acknowledgements.
We are indebted to G. Gibbons for his insights and advices. We benefitted from discussions and correspondence also with P. C. Aichelburg, T. Damour, L. Diósi, M. Elbistan, S. Kar, B. Kocsis, V. Komornik, K. Mitman and S. Silagadze. JB is grateful to SYSU for hospitality extended to him in Zhuhai. PAH thanks the Erwin Schrödinger Institute (ESI, Vienna) for hospitality during the Workshop Carrollian Physics and Holography-CDFG-2024. The work of JB has been supported in part by the NKFIH grant K134946. PMZ was partially supported by the National Natural Science Foundation of China (Grant No. 12375084).

### v6

@@ V Higher-derivative profiles
Similar results can be obtained for the higher-derivative profiles proposed in GibbHaw71 ; LongMemory ; ZZHflyby .
@@
This curious “half-sidedness” is explained as follows. Assuming that $k>0$, DM is obtained for $X^{2}$ by fine-tuning the amplitude, $k=k_{crit}^{(2)}$. This does not work for $X^{1}$, though, unless $X^{1}\equiv 0$, as confirmed by Figs.11 and FIG.12.
Figure 12: Fine-tuning one of the sectors for DM yields VM for the other, when the latter trajectory is identically zero.
However we could proceed also with the opposite casting and fine-tune the amplitude for $X^{1}$ to get $k=k_{crit}^{(1)}$. But the profiles have opposite signs, therefore
@@
Models with higher-order derivatives of the Gaussian follow a similar pattern. For the 4th order derivative profile, for example, we would get again half-DM effect, etc. We now explain the reason by studying the behavior under reversing the sign of $U$,
$U\to-U\,.$ (V.5)
We start with an even seed profile,
${{\mathcal{A}}}^{0}(-U)={{\mathcal{A}}}^{0}(U)\,$ (V.6)
@@
The sign-reversal (V.5) leaves invariant the 2nd-order $U$-derivative terms in (I.3) but has an $\ell$-dependent effect on the equations of motion :
@@
implying that the two transverse components in (I.3a)-(I.3b) are merely interchanged under (V.5) (as noticed for flyby in sect.III). Thus they satisfy the same pair of equations up to permutation. Therefore the DM condition holds either for both or for none of them. When it does, then “full DM” is obtained for both coordinates as illustrated in FIGs. 7-8 for $\ell=1$, and in FIG.14 for $\ell=3$.
$U$-reversal symmetry (V.5) allows us to generalize an observation made in sect.IV: for the chosen normalization the $U\to-U$ symmetry implies
$X^{1}(\infty)\,X^{2}(\infty)=1\,.$ (V.9)
Thus the larger one of the components is, the smaller the other will be. A higher amplitude improves the precision of DM.
@@
Each of the transverse equations (I.3a) and (I.3b) is left invariant under (V.5)
@@ VI Longitudinal motion
$\mathfrak{m}=M\,,$ (VI.4)
then, setting
$Z=V+\frac{1}{2}U$ (VI.5)
we get,
$Z(U)=V_{0}=\mathop{\rm const.}\nolimits$ (VI.6)
i.e.,  we get DM also in the longitudinal $Z$-direction, see FIG.#(14) of DM-1 .
@@ VII Conclusion
Test particles hit by a sandwich gravitational wave exhibit, generically, the velocity effect (VM) Ehlers ; Sou73 ; AiBalasin  : after the wave has passed, the particles fly apart with diverging constant, non-zero velocity ShortMemory ; LongMemory . Zel’dovich and Polnarev suggested instead that flyby would approximately generate pure displacement (DM) with vanishing relative velocity ZelPol . Our main result is to confirm and indeed refine the statement of Zel’dovich and Polnarev. Both numerical and analytical evidence show that a judicious choice of the wave parameters yields pure displacement DM-1
We focus our attention at linearly polarized gravitational waves whose profile is a derivative of order =0, $\ell=0,\,1,\ \dots$ of a seed profile ${\mathcal{A}}^{0}$, (V.7). The latter can be, for example, a Gaussian GibbHaw71 or its analytical Doppelgänger, the Pöschl-Teller potential PTeller ; Chakra ; DM-1 ; ZZHflyby . Other promising candidates (not studied in the paper) could be Rosen-Morse type Scarf or inverse-power AndrPrenc ; Ilderton ; ZhaoU-4 profiles. Analytically solvable models can be built also from square-profile Chakraborty:2022qvv ; ZZHflyby .
@@
Preliminary versions of this research ZZHflyby were presented by PAH at the Wigner Institute in Budapest (April 5 2024), at the Workshop Carrollian Physics and Holography-CDFG-2024 organised at the Schrödinger Institute in Vienna (April 17 2024), and at the conference “Conformal anomalies: theory and applications 2024, https://indico.math.cnrs.fr/event/10718/ Tours (May 7 2024). P-M Zhang presented related results in ZZHflyby .
Related ideas were put forward also in Bondi57 ; Sou73 ; AiBalasin ; GriPol . Recent results on the Memory Effect and the DM or VM puzzle include refs. Chakraborty:2022qvv ; DM-1 ; Jibril ; ZZHflyby ; Mitman ; HarteOancea ; Arpan ; DeyKar .
@@ Acknowledgements.
We are indebted to G. Gibbons for his insights and advices. We benefitted from discussions and correspondence also with P. C. Aichelburg, T. Damour, L. Diósi, M. Elbistan, S. Kar, B. Kocsis, V. Komornik, K. Mitman and S. Silagadze. JB is grateful to SYSU for hospitality extended to him in Zhuhai. PAH thanks the Erwin Schrödinger Institute (ESI, Vienna) for hospitality during the Workshop Carrollian Physics and Holography-CDFG-2024.